Solving Problem Using Linear System
18. A train leaves Toronto for Montreal at the same time as another train leaves Montreal for Toronto. The cities are 500 km apart. The trains pass each other 2 h later. The train from Montreal is travelling 50 km/h faster than the one from Toronto do the trains pass each other?
Can you please help me with this question?
Thanks a lot!!!

Answer by ankor@dixie-net.com(22740)

8. A train leaves Toronto for Montreal at the same time as another train leaves
Montreal for Toronto. The cities are 500 km apart. The trains pass each other
2 h later. The train from Montreal is traveling 50 km/h faster than the one
from Toronto. Assume you mean,"How far from Toronto do the trains pass each other?"
:
Let s = the speed of the train from Toronto
then
(s+50) = the speed of the train from Montreal
:
When the two trains meet, they will have traveled a total of 500 mi
Write a distance equation; dist = time * speed
2(s+50) + 2s = 500
2s + 100 + 2s = 500
2s + 2s = 500 - 100
4s = 400
s = 100 km/hr is the Toronto train
in 2 hrs the the Toronto train will be 200 km from Toronto
and
check it by finding that 2(150) = 300 km is the Montreal train from Montreal